# Fine-Grained and Multi-Dimensional Metrics for Document-Level Machine Translation

### v3

@@ Abstract
Large language models (LLMs) have excelled in various NLP tasks, including machine translation (MT), yet most studies focus on sentence-level translation. This work investigates the inherent capability of instruction-tuned LLMs for document-level translation (docMT). Unlike prior approaches that require specialized techniques, we evaluate LLMs by directly prompting them to translate entire documents in a single pass. Our results show that this method improves translation quality compared to translating sentences separately, even without document-level fine-tuning. However, this advantage is not reflected in BLEU scores, which often favor sentence-based translations. We propose using the LLM-as-a-judge paradigm for evaluation, where GPT-4 is used to assess document coherence, accuracy, and fluency in a more nuanced way than n-gram-based metrics. Overall, our work demonstrates that instruction-tuned LLMs can effectively leverage document context for translation. However, we caution against using BLEU scores for evaluating docMT, as they often provide misleading outcomes, failing to capture the quality of document-level translation  [^1] Our code and the outputs from GPT4-as-a-judge are available at https://github.com/EIT-NLP/BLEUless_DocMT.
@@ 1 Introduction
  • We propose using the LLM-as-a-judge paradigm with tailored prompts for more accurate docMT evaluation.
@@
  • We recommend against using BLEU scores for docMT, as they fail to capture discourse-level phenomena and can often provide misleading results.
@@ Results.
While some judgments by the LLM-as-a-judge may appear reasonable, certain nuances may still be misinterpreted due to unique human perspectives. To validate the alignment between our LLM-as-a-judge paradigm and human evaluations, we conducted experiments to assess agreement. For each model in both ST3 and DOC, we used 10 samples per translation direction and asked human evaluators to respond with a simple ‘yes’ or ‘no’ regarding their agreement with the LLM-as-a-judge’s assessments according to our metrics.
@@
On the right side of first case in Box 2, the translation of “Hunan” remains consistent throughout the document, illustrating the LLM’s capability to leverage context and capture inter-sentence dependencies. Conversely, on the left side, we see an erroneous translation where “Hunan” is rendered as "Lake South” in “Lake South Red Tea” and simply as “Lake” in “Lake Red Source”. Notably, the model in ST3 correctly translates “Hunan” in other parts of the text. In this case, although ST3 achieves a BLEU score approximately 11.88 points higher than DOC, it is evident that DOC provides more coherent wording and aligns better with natural human expression.
@@ 5 Conclusion
In this work, we investigate the performance of instruction-tuned LLMs in document-level machine translation (docMT), comparing the translation of entire documents in a single pass to translating individual sentences that are then concatenated. Our findings show that translating entire documents yields better results, as the model can capture inter-sentence dependencies and maintain discourse coherence, even without explicit fine-tuning for docMT tasks. However, evaluating these improvements is challenging. Traditional metrics like BLEU fail to consider discourse-level phenomena, often favoring sentence-level translations and producing misleading results. To address this limitation, we propose the LLM-as-a-judge approach, utilizing GPT-4 to assess specific aspects of discourse through tailored prompts. This method enhances interpretability and can be adapted for evaluating translation quality in other domains.

### v4

@@ Abstract
Large language models (LLMs) have excelled in various NLP tasks, including machine translation (MT), yet most studies focus on sentence-level translation. This work investigates the inherent capability of instruction-tuned LLMs for document-level translation (docMT). Unlike prior approaches that require specialized techniques, we evaluate LLMs by directly prompting them to translate entire documents in a single pass. Our results show that this method improves translation quality compared to translating sentences separately, even without document-level fine-tuning. However, this advantage is not reflected in BLEU scores, which often favor sentence-based translations. We propose using the LLM-as-a-judge paradigm for evaluation, where GPT-4 is used to assess document coherence, accuracy, and fluency in a more nuanced way than n-gram-based metrics. Overall, our work demonstrates that instruction-tuned LLMs can effectively leverage document context for translation. However, we caution against using BLEU scores for evaluating docMT, as they often provide misleading outcomes, failing to capture the quality of document-level translation. [^1] Our code and the outputs from GPT4-as-a-judge are available at https://github.com/EIT-NLP/BLEUless_DocMT
@@ 1 Introduction
  • We propose using the LLM-as-a-judge paradigm with multiple prompts that assess different aspects of translated text to achieve a more targeted and accurate evaluation.
  • We recommend against using d-BLEU scores for docMT, as they fail to capture discourse-level phenomena and can often provide misleading results.
@@ Results.
While some judgments by the LLM-as-a-judge may appear reasonable, certain nuances may still be misinterpreted due to unique human perspectives. To validate the alignment between our LLM-as-a-judge paradigm and human evaluations, we conducted experiments to assess agreement. For each model in both ST3 and DOC, we used 10 samples per translation direction and asked human evaluators to respond with a simple “yes” or “no” regarding their agreement with the LLM-as-a-judge’s assessments according to our metrics.
@@
On the right side of first case in Box 2, the translation of “Hunan” remains consistent throughout the document, illustrating the LLM’s capability to leverage context and capture inter-sentence dependencies. Conversely, on the left side, we see an erroneous translation where “Hunan” is rendered as “Lake South” in “Lake South Red Tea” and simply as “Lake” in “Lake Red Source”. Notably, the model in ST3 correctly translates “Hunan” in other parts of the text. In this case, although ST3 achieves a BLEU score approximately 11.88 points higher than DOC, it is evident that DOC provides more coherent wording and aligns better with natural human expression.
@@ 5 Conclusion
In this work, we investigate the performance of instruction-tuned LLMs in document-level machine translation (docMT), comparing the translation of entire documents in a single pass to translating individual sentences that are then concatenated. Our findings show that translating entire documents yields better results, as the model can capture inter-sentence dependencies and maintain discourse coherence, even without explicit fine-tuning for docMT tasks. However, evaluating these improvements is challenging. Traditional metrics like d-BLEU fail to consider discourse-level phenomena, often favoring sentence-level translations and producing misleading results. To address this limitation, we propose the LLM-as-a-judge approach, utilizing GPT-4 to assess specific aspects of discourse through tailored prompts. This method enhances interpretability and can be adapted for evaluating translation quality in other domains.